Why is the relevant range key between variable and marginal costs?
Answer #1

Total Cost=TC =F+V*Q

F=Fixed Cost

V=Variable Cost/unit

Q=Units of production

Marginal cost is the cost of production of one additional unit

Variable cost is proportional to the units produced.

Since fixed cost is constant, the total cost TC will increase by V for one additional unit of production.

Hence ,so long Fixed cost is constant, Marginal Cost=Variable cost per unit=V.

For example,

If fixed cost is $100,000(for the range of production of 500 to 1000 units)

Variable Cost per unit=$100

Cost of producing 500 units=$100,000+$100*500=$150,000

Cost of producing 501 units=$100,000+$100*501=$150,100

Marginal cost =150100-150000=$100 =Variable cost per unit/

In this case relevant range of Fixed Cost is 500 to 1000 units.

If quantity of production Q is more than 1,000 units,the Fixed cost changes due to addition of machines etc.

Assume it increases to $150,000

TC for Q=1000

TC=$100,000+$100*1000=$200,000

For Q=1001,

TC=$150,000+$100*1001=$250,100

In this case ,Marginal cost =$250,100-$200,000=$50,100

Hence Marginal Cost is much higher than the variable cost

Hence, relevant range is key between variable and marginal costs.

Only within the relevant range, marginal cost equals the variable cost

The relevant range is the activity level that is bounded by a minimum and maximum amount. Within relevant range, certain cost level is fixed. Outside of that relevant range, the cost level will change.